OpenStudy (anonymous):

You have seen that the total number of lights in a triangular lighting rig is given by T(n) = n2/2 + n/2, while T(n+1) = n2/2 + 3n/2 + 1. If T(n) times T(n+1) = f(n), what is the reduced form of f(n)? f(n)=n^4/4+n^3+5n^2/4+n/2?? is this correct

OpenStudy (asnaseer):

yes - looks right to me

OpenStudy (asnaseer):

depends what they mean by "reduced form" as it could aso be written as:\[\frac{n}{4}(n+1)(n^2+3n+2)\]

OpenStudy (asnaseer):

and further simplified to:\[\frac{n}{4}(n+1)(n+1)(n+2)=\frac{n}{4}(n+1)^2(n+2)\]